OpenStudy (anonymous):

What values of x satisfies log2(3x) + log2(2x) = 3?

OpenStudy (mrnood):

log a + log b = log (a*b) so in this case log (3x) + log (2x) = log (6x^2) = 3

OpenStudy (anonymous):

How do I deal with the log base 2

OpenStudy (mrnood):

base of log is irrelevent: logn (x) = m means n^m = x

OpenStudy (anonymous):

Ahhhhhh I got it

OpenStudy (zenmo):

Okay, similar to @MrNood \[\log_{2} (3x)+\log_{2} (2x)=3\] Using the identity: loga+logb=log(a*b) \[\log_{2} (6x^2)=3\]

OpenStudy (zenmo):

good job then. :)

OpenStudy (anonymous):

I just forgot all the identities, that's all

OpenStudy (anonymous):

log2(6x^2) = 3 2^3 = 6x^2 4/3 = x^2

OpenStudy (anonymous):

Mmm I see, thanks

OpenStudy (zenmo):

yep, do you know how to simplify ur answer?

OpenStudy (zenmo):

unless that answer format is accepted

OpenStudy (anonymous):

Yea x = (4/3)^(1/2)